# Attention as an Adaptive Filter

### v2

@@ 5 Radial-Tangential Model
5.1 Itô SDE with Radial–Tangential Dynamics and Noise
Consider the linear stochastic differential equation:
@@ 5.2 Propagation of Uncertainty through the RT-SDE
5.3 Radial–Tangential Kalman IRLS
@@ 5.4 Algorithm
Complex linear layers: $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$
@@
Lift inputs to complex form: $\mathbf{Z}=[\mathbf{Z}_{r},0]^{\top}$
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z}))$
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}}))$
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}}))$
@@
$\mathbf{V}^{C}_{r}[i,j]=\kappa(|i-j|)\sigma_{r}^{2}+e^{-2\alpha^{2}\Delta t|i-j|}\eta_{r}^{2}$
$\mathbf{V}^{C}_{t}[i,j]=\kappa(|i-j|)\sigma_{t}^{2}+e^{-2\alpha^{2}\Delta t|i-j|}\eta_{t}^{2}$
@@
Dynamic precision kernels:
$\mathbf{P}^{C}_{\text{rad}}[i,j]=1/\sigma_{V_{r}}^{2}(|t_{i}-t_{j}|),\quad\mathbf{P}^{C}_{\text{tan}}[i,j]=1/\sigma_{V_{t}}^{2}(|t_{i}-t_{j}|)$
@@
$\mathbf{\tilde{A}}_{\text{rad}}[i,j]=-\log(\mathbf{V}^{C}_{\text{rad}}[i,j]+\mathbf{R}_{\text{rad}}^{2}[i,j]+\epsilon),$
$\mathbf{\tilde{A}}_{\text{tan}}[i,j]=-\log(\mathbf{V}^{C}_{\text{tan}}[i,j]+\mathbf{R}_{\text{tan}}^{2}[i,j]+\epsilon)$
@@
$\big(\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big)\leftarrow\mathcal{L}_{P}(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+}))$
@@ 5.5 Interpreting the Transformer as an MLE on a Hypersphere
For small angles, the attention update is largely aligned with the tangent space and any component along $u$ is partially corrected by the subsequent geodesic normalization, so we omit the explicit tangent projection for simplicity.
@@
Parameters: AFA layer $\mathrm{AFA}_{\mathrm{dir}}$, scalar AFA layer $\mathrm{AFA}_{\mathrm{mag}}$, output complex linear layer $\mathcal{L}_{O}$, real-valued FFN layers: $\mathrm{FFN}=\mathcal{L}_{2}\circ\mathrm{ReLU}\circ\mathcal{L}_{1}$, stability const. $\epsilon$.
Lift real inputs to complex by concatenating zeros: $\mathbf{Z}=\left[\mathbf{Z}_{r},\mathbf{0}\right]^{\top}\in\mathbb{C}^{2\times d\times m}$
Complex projections:
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big).$
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big)$,
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big)$.
@@
Output complex linear projection:
$(\operatorname{Re}(\mathbf{Z_{o}}),\operatorname{Im}(\mathbf{Z_{o}}))\leftarrow\mathcal{L}_{O}(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z}))$
@@
$\mathbf{Z}^{\mathrm{out}}=\mathbf{Z}_{r}+\mathrm{FFN}(\operatorname{Re}(\mathbf{Z_{o}}))$
@@ 6.4 Neural State Space Models
The HiPPO framework improves memory retention in SSMs by introducing recurrent memory with optimal polynomial projections, which allows the model to compress and retain important information from the entire history in a structured latent state space (Gu et al., 2020). The Structured State Space Sequence model (S4) designs efficient state space layers that scale linearly with sequence length, using structured representations of the system matrices (Gu et al., 2022). The Toeplitz Neural Network (TNN) efficiently models sequences using relative position–encoded Toeplitz matrices and a lightweight Relative Position Encoder, enabling fast token mixing and extrapolation to much longer sequences with log-linear complexity (Qin et al., 2023).
@@
Both S4 and our model exploit a diagonal-plus-low-rank (DPLR) structure to reduce the cost of inverting a dense $d\times d$ matrix, using the Sherman-Morrison formula for efficiency: a diagonal part captures independent modes, and a rank-1 correction captures global interactions. In S4, the DPLR matrix models deterministic state dynamics, determining how past inputs influence future states. In our model, the DPLR structure appears in the noise covariance matrices in a stochastic differential equation (SDE) formulation, which allows us to represent how uncertainty propagates through the dynamics.
6.5 Complex-Valued Neural Networks
Trabelsi et al. (2018) introduced practical components for training deep complex-valued neural networks—complex convolutions, batch normalization, weight initialization, and activation functions—and demonstrated their effectiveness on vision and audio tasks, showing competitive performance with real-valued networks (Trabelsi et al., 2018). Eilers and Jiang (2023) developed core components for a complex-valued transformer, including complex Scaled Dot-Product Attention and layer normalization, enabling fully complex-valued models that match real-valued performance while improving robustness to overfitting on complex-valued tasks (Eilers and Jiang, 2023). Ran-Milo et al. (2024) show that complex-valued SSMs can express oscillatory and long-range mappings with moderate dimensions and parameters, while real SSMs require much larger or exponentially scaled parameters for the same expressivity (Ran-Milo et al., 2024). Leng et al. develop a framework for complex-valued neural networks and transformers in wireless communications, showing that CVNNs require fewer layers than real-valued networks and introducing complex-valued embedding, encoding, decoding, and output modules (Leng et al., 2025).
@@ 6.7 Structuring Attention with Priors and Positional Biases
Some recent works focus on understanding the mathematical structure of rotary embeddings and extending them to higher dimensions or more flexible parameterizations. Yu et al. introduce ComRoPE, which replaces fixed 2D rotation matrices with trainable commuting angle matrices (Yu et al., 2025). Liu et al. use commuting skew-symmetric generators to define an $N$-dimensional RoPE, generalizing 1D and 2D forms and allowing learnable inter-dimensional interactions (Liu et al., 2025). CoPE (Amballa, 2025), and ComplexFormer (Shao et al., 2025) extend RoPE into the complex domain. Wu et al. observe that the same rotation trick used in Transformers (on queries and keys) can be applied to the matrices in an SSM formulation (Wu et al., 2025).
@@
Several papers have explored element-wise (in the embedding dimension) attention mechanisms. Elliptical Attention replaces the traditional dot-product attention with a Mahalanobis distance-based mechanism (Nielsen et al., 2024). McCarter investigates replacing scaled dot-product attention with negative-log Euclidean distance (McCarter, 2023). Feng uses an exponentiated Euclidean distance as a similarity metric (Feng, 2025). Bacochina et al. replace matrix multiplications with learned spatial weight arrays applied directly via element-wise multiplication, focusing attention along rows and columns (Bacochina and de Souza, 2023). Wu et al. introduce the DA-Transformer, which reweights attention scores using learnable functions of token distances (Wu et al., 2021)
@@ 6.9 Modeling Uncertainty in Neural Networks
6.10 Structured Low-Rank Covariances for Efficient Propagation
Several methods have been developed to efficiently represent and propagate high-dimensional covariance matrices using low-rank plus diagonal approximations in Riccati-like equations. Pnevmatikakis et al. propose a low-rank perturbative approach for Kalman Filtering and smoothing, approximating the state covariance as a small-rank update to the prior (Pnevmatikakis et al., 2014). Saibaba et al. represent the covariance as a weighted sum of the process noise and a low-rank term from a generalized eigenvalue problem (Saibaba et al., 2015). Bonnabel et al. approximate large, time-dependent PSD matrices with low-rank plus diagonal structures, allowing linear-time storage, inversion, and integration in high-dimensional Riccati flows (Bonnabel et al., 2024). Schmidt et al. (2021) introduce the Rank-Reduced Kalman Filter, projecting the Lyapunov equations onto a low-rank manifold and updating only the covariance column space, reducing complexity from cubic to quadratic (or linear) while recovering the exact filter in the full-rank limit and outperforming ensemble-based approaches (Schmidt et al., 2024).
@@ 6.12 Geometric Views of Transformers
Geshkovski et al. model Transformers as interacting particle systems evolving on the unit sphere, and find that self-attention induces nonlinear dynamics leading to clustering of token embeddings (Zhang et al., 2021; Geshkovski et al., 2024). Molina interprets Transformers geometrically, showing that LayerNorm constrains hidden states to a hypersphere and attention moves “word particles” along this surface (Molina, 2023). Brody et al. show that LayerNorm shapes Transformer hidden states geometrically — projecting them onto a hyperplane and scaling them — which enhances the expressiveness of attention (Brody et al., 2023). Loshichilov et al. introduce nGPT, a Transformer variant that normalizes all embeddings and network weights on the hypersphere, enabling cosine-similarity-based updates with learnable per-dimension step sizes, resulting in faster convergence and more stable training (Loshchilov et al., 2025). Gupta et al. observe that LayerNorm can be interpreted geometrically as projecting a vector orthogonally to the uniform vector, normalizing it, and scaling by the dimensionality, making it irreversible and largely redundant compared to RMSNorm, which achieves similar hidden representations more efficiently (Gupta et al., 2025).

### v3

@@ 5 Radial-Tangential Model
We choose the radial direction for anisotropic noise because it is the only direction that co-rotates perfectly with the angular dynamics of the state vector. This co-rotation causes the time-dependent rotation terms to cancel out in the covariance integral, enabling us to express fixed anisotropic noise in the token frame, while maintaining a closed-form expression for the propagated covariance.
5.1 Radial-Tangential SDE
Consider the linear Itô stochastic differential equation:
@@ 5.2 Propagation of Uncertainty through the RT-SDE
5.3 Radial–Tangential IRLS
@@
As before, we add additional measurement noise at point $i$, $\gamma_{r}^{2},\gamma_{t}^{2}$, and let:
$\sigma_{Vr,ij}^{2}\leftarrow\sigma_{Vr,ij}^{2}+\gamma_{r}^{2},\quad\sigma_{Vt,ij}^{2}\leftarrow\sigma_{Vt,ij}^{2}+\gamma_{t}^{2},$
@@ 5.4 Algorithm
Linear layers: $\mathcal{L}_{Q}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{K}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{V}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{P}^{\mathbb{C}\rightarrow\mathbb{R}}$;
@@
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big),$
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big),$
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big).$
@@
$\mathbf{V}^{Z}_{\text{rad}}[i,j]=\kappa(|i-j|)\sigma_{r}^{2}+\eta_{r}^{2}e^{-2\alpha^{2}\Delta t|i-j|}+\gamma_{r}^{2}$
$\mathbf{V}^{Z}_{\text{tan}}[i,j]=\kappa(|i-j|)\sigma_{t}^{2}+\eta_{t}^{2}e^{-2\alpha^{2}\Delta t|i-j|}+\gamma_{t}^{2}$
@@
$\mathbf{\tilde{A}}_{\text{rad}}[i,j]=-\log(\mathbf{V}^{Z}_{\text{rad}}[i,j]+\mathbf{R}_{\text{rad}}^{2}[i,j]+\epsilon),$
$\mathbf{\tilde{A}}_{\text{tan}}[i,j]=-\log(\mathbf{V}^{Z}_{\text{tan}}[i,j]+\mathbf{R}_{\text{tan}}^{2}[i,j]+\epsilon)$
@@
$\operatorname{Re}(\mathbf{Z_{p}})\leftarrow\mathcal{L}_{P}^{\mathbb{C}\rightarrow\mathbb{R}}\big(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+})\big)$
@@ 5.5 Interpreting the Transformer as an MLE on a Hypersphere
For small angles, the attention update is largely aligned with the tangent space and any component along $\mathbf{u}_{\mathbf{zs}i}^{(k)}$ is partially corrected by the subsequent geodesic normalization, so we omit the explicit tangent projection for simplicity.
@@
Parameters: AFA layers: $\mathrm{AFA}_{\mathrm{dir}}$, $\mathrm{AFA}_{\mathrm{mag}}$;
Linear layers: $\mathcal{L}_{Q}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{K}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{V}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{P}^{\mathbb{C}\rightarrow\mathbb{R}}$; FFN: $\mathrm{FFN}=\mathcal{L}_{2}\circ\mathrm{ReLU}\circ\mathcal{L}_{1}$, stability const. $\epsilon$.
Complex-valued Projections:
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big),$
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big),$
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big).$
@@
Output projection:
$\operatorname{Re}(\mathbf{Z_{p}})\leftarrow\mathcal{L}_{P}^{\mathbb{C}\rightarrow\mathbb{R}}\big(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+})\big)$
@@
$\mathbf{Z}^{\mathrm{out}}=\mathbf{Z}_{r}+\mathrm{FFN}(\operatorname{Re}(\mathbf{Z_{p}}))$
@@ 6.3 Neural State Space Models
The HiPPO framework improves memory retention in SSMs by introducing recurrent memory with optimal polynomial projections, which allows the model to compress and retain important information from the entire history in a structured latent state space (Gu et al., 2020). The Structured State Space Sequence model (S4) designs efficient state space layers that scale linearly with sequence length, using structured representations of the system matrices (Gu et al., 2022). S4 exploits a diagonal-plus-low-rank (DPLR) structure to reduce the cost of inverting a dense $d\times d$ matrix, using the Sherman-Morrison formula for efficiency: a diagonal part captures independent modes, and a rank-1 correction captures global interactions. The Toeplitz Neural Network (TNN) efficiently models sequences using relative position–encoded Toeplitz matrices and a lightweight Relative Position Encoder, enabling fast token mixing and extrapolation to much longer sequences with log-linear complexity (Qin et al., 2023).
@@
Like S4, AFA makes use of a diagonalizable linear dynamics model to simplify the computation. In S4, a DPLR matrix models deterministic state dynamics. In our model, a DPLR structure appears in the noise covariance matrices in a stochastic differential equation (SDE) formulation, which allows us to represent how uncertainty propagates through the dynamics.
@@ 6.4 Attention as a Structured State Space Computation
6.5 Complex-Valued Neural Networks
Trabelsi et al. (2018) introduce practical components for training deep complex-valued neural networks—complex convolutions, batch normalization, weight initialization, and activation functions—and demonstrated their effectiveness on vision and audio tasks, showing competitive performance with real-valued networks (Trabelsi et al., 2018). Eilers and Jiang (2023) develop core components for a complex-valued Transformer, including CAtt — a symmetric, rotationally invariant feature similarity kernel in the complex plane — and complex-valued layer normalization (Eilers and Jiang, 2023). Ran-Milo et al. (2024) show that complex-valued SSMs can express oscillatory and long-range mappings with moderate dimensions and parameters, while real SSMs require much larger or exponentially scaled parameters for the same expressivity (Ran-Milo et al., 2024). Leng et al. develop a framework for complex-valued neural networks and Transformers in wireless communications, showing that CVNNs require fewer layers than real-valued networks, and introducing complex-valued embedding, encoding, decoding, and output modules (Leng et al., 2025).
@@ 6.6 Structuring Attention with Priors and Positional Biases
Some recent works focus on understanding the mathematical structure of rotary embeddings and extending them to higher dimensions or more flexible parameterizations. Yu et al. introduce ComRoPE, which replaces fixed 2D rotation matrices with trainable commuting angle matrices (Yu et al., 2025). Liu et al. use commuting skew-symmetric generators to define an $N$-dimensional RoPE, generalizing 1D and 2D forms and allowing learnable inter-dimensional interactions (Liu et al., 2025). Ostmeier et al. introduce LieRE (Lie Rotational Positional Encodings), which generalizes RoPE by accommodating $n$-dimensional inputs (Ostmeier et al., 2025). CoPE (Amballa, 2025), and ComplexFormer (Shao et al., 2025) extend RoPE into the complex domain. Wu et al. observe that the same rotation trick used in Transformers (on queries and keys) can be applied to the matrices in an SSM formulation (Wu et al., 2025).
@@
Several papers have explored element-wise (in the embedding dimension) attention mechanisms. Elliptical Attention replaces the traditional dot-product attention with a Mahalanobis distance-based mechanism (Nielsen et al., 2024). McCarter investigates replacing scaled dot-product attention with negative-log Euclidean distance (McCarter, 2023). Feng uses an exponentiated Euclidean distance as a similarity metric (Feng, 2025). Bacochina et al. replace matrix multiplications with learned spatial weight arrays applied directly via element-wise multiplication, focusing attention along rows and columns (Bacochina and de Souza, 2023). Wu et al. introduce the DA-Transformer, which reweights attention scores using learnable functions of token distances (Wu et al., 2021). A 2024 survey by Zhao et al. systematically categorizes existing Positional Encodings (PEs) and related methods, focusing on their role in enabling length extrapolation of Transformers, while highlighting the need for a solid theoretical foundation to fully understand and achieve long-context generalization (Zhao et al., 2024).
6.7 Comparison to Rotary Positional Encodings
As we noted previously, Isotropic AFA leads naturally to a complex-valued generalization of dot product attention $\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}})$, where the imaginary part of the state transition matrix $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}$ is a complex-valued generalization of a standard rotational matrix. Its action, which involves element-wise complex multiplication (rotation) on the queries and keys, provides a relative complex positional encoding structure that encompasses the geometry of RoPE (Su et al., 2023). The assumption of isotropic decay $\mathbf{A}=\alpha\mathbf{I}+\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, which ensures that the system’s state propagation neatly decouples into a uniform exponential decay $\mathbf{E_{qk}}$ and a complex rotation $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}$, is the same as that used in xPos (Sun et al., 2022) and the Retentive Network (RetNet) (Sun et al., 2023).
[…]
By contrast, RoPE (and xPos) remain entirely real-valued except for a fixed $2\times 2$ block rotation. The input $\mathbf{X}\in\mathbb{R}^{d\times m}$ is projected to $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{d\times m}$ using standard real linear layers, then rotated as $\mathbf{Q}^{\prime}=\mathbf{R_{q}}\mathbf{Q}$, $\mathbf{K}^{\prime}=\mathbf{R_{k}}\mathbf{K}$. The attention score is computed with a standard real inner product $\mathbf{Q}^{\prime}(\mathbf{K}^{\prime})^{\top}$. No complex weights are learned, and the imaginary components serve only to implement the positional rotation.
The distinction becomes clear in the attention scores:
$\text{Score}_{\text{AFA}}=\mathrm{Re}\!\left(\mathbf{\tilde{q}}^{\dagger}\mathbf{\tilde{k}}\right)=\mathbf{\tilde{q}}_{r}^{\top}\mathbf{\tilde{k}}_{r}+\mathbf{\tilde{q}}_{i}^{\top}\mathbf{\tilde{k}}_{i},\qquad\text{Score}_{\text{RoPE}}\;\propto\;(\mathbf{R}_{m}\mathbf{q})(\mathbf{R}_{n}\mathbf{k})=\mathbf{\tilde{q}}_{r}^{\top}\mathbf{\tilde{k}}_{r}.$
RoPE discards the contribution of the imaginary parts $(\mathbf{\tilde{q}}_{i},\mathbf{\tilde{k}}_{i})$, whereas AFA fully exploits the complex structure to couple the real and imaginary components.
Additionally, AFA integrates uncertainty priors/biases into the attention weights using propagated covariance kernels. Furthermore, AFA rotates the values to the rotated query-key frame before multiplying by the attention matrix, whereas RoPE does not, keeping its values static. After computing the attention scores, AFA applies an inverse rotation $(\bm{\tilde{\Phi}}_{\mathbf{v}}^{+})^{\dagger}$ that maps the values back into the original frame:
Finally, AFA applies a prediction step using the same state transition matrix used for estimation:
$\mathbf{Z}_{\mathbf{v}}^{+}=e^{\alpha\Delta t}e^{\bm{\omega}\Delta t}\odot\mathbf{\bar{Z}_{v}}$
@@ 6.11 Geometric Views of Transformers
Geshkovski et al. model Transformers as interacting particle systems evolving on the unit sphere, and find that self-attention induces nonlinear dynamics leading to clustering of token embeddings (Zhang et al., 2021; Geshkovski et al., 2024). Molina interprets Transformers geometrically, showing that LayerNorm constrains hidden states to a hypersphere and attention moves “word particles” along this surface (Molina, 2023). Brody et al. show that LayerNorm shapes Transformer hidden states geometrically — projecting them onto a hyperplane and scaling them — which enhances the expressiveness of attention (Brody et al., 2023). Loshichilov et al. introduce nGPT, a Transformer variant that normalizes all embeddings and network weights on the hypersphere, enabling cosine-similarity-based updates with learnable per-dimension step sizes, resulting in faster convergence and more stable training (Loshchilov et al., 2025). Gupta et al. observe that LayerNorm can be interpreted geometrically as projecting a vector orthogonally to the uniform vector, normalizing it, and scaling by the dimensionality, making it irreversible and largely redundant compared to RMSNorm, which achieves similar hidden representations more efficiently (Gupta et al., 2025). Post-projection normalization is explored in prior work such as Query-Key Normalization for Transformers (Henry et al., 2020), though this work does not connect this choice to geometrical considerations.